I need help finding the %M/V

The protein albumen, found in human blood, has a m
Answer #1

Converting units

% m/v = 40 g/L = 40% m/v

ppm = mg/L

       = 40 x 1000/L

       = 40,000 ppm

M = 40 g/66,500 x L

    = 0.0006 M